# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v1

@@ 5 Experiments
In this section, we describe our experiments based on synthetic and real-world data.
@@
We consider three synthetic setups with different choices of priors in $P$. For the first setup, the priors have varying kernels selected according to one of the following kernels: i) squared-exponential kernel, Matérn kernel with $\nu=5/2$, ii) Matérn kernel with $\nu=3/2$, iii) periodic kernel with period $\rho=5$, iv) linear kernel with $v=0.05^{2}$, and, v) the rational quadratic kernel with $\alpha=0.5$. All kernels use a lengthscale of $1.0$ and are scaled s.t. $k(x,\tilde{x})\leq 1$. In addition, the mean function for all priors are zero everywhere. For the second setup, the priors use the squared-exponential kernel with lengthscales $4$, $2$, $1$ or $1/2$. The third setup is based on the toy problem in an earlier version of Ziomek et al. (2024b) where the mean functions vary. The first mean function has 10 smaller ‘hills’ across the first half of the interval. The four other mean functions are the same except the height of one of the hills is doubled in height, see Fig. 3. All priors in the third setup use a squared-exponential kernel with lengthscale 1. For all of the setups, the true prior $p^{*}$ is sampled uniformly from $P$, 500 arms are equidistantly spaced in $[0,20]$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$. The prior elimination methods use $\delta=0.05$. All models are evaluated on 500 seeds on each setup. As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$. Regardless of the selected prior, HP-GP-TS observes an observation $y_{t}$ to update the hyperposterior $P_{t}$. Hence, greedily selecting the prior could reduce unnecessary exploration. In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{0}=\{p^{*}\}$.
Figure 6: Average entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS. The dashed reference values correspond to entropies of discrete distributions with prob. $q$ on one choice and prob. $\frac{1-q}{|P|-1}$ on the other $|P|-1$ choices.
@@
The cumulative regret for the three synthetic experiments is shown in Fig. 4. Across all three experiments, we observe that HP-GP-TS has lower regret than the other methods and performs close to the oracle GP-TS. For the kernel and hill experiments, PE-GP-TS has lower regret than the oracle GP-UCB. Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods. MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and hill experiments but has significantly higher regret for the kernel experiment. The greedy selection causes under-exploration for MAP GP-TS in certain instances.
In the three left-most columns of Fig. 5, the average amount of priors remaining is shown over time for the synthetic experiments (for the PE-methods methods). For the hill experiment, no priors are eliminated for either PE-GP-TS nor PE-GP-UCB. For the kernel and lengthscale experiment, PE-GP-TS eliminates more priors than its UCB-counterpart. However, note that, on average, less than 0.15 priors are eliminated in the lengthscale experiment and slightly more than 1.2 priors in the kernel experiment.
Similarly, in Fig. 6, the average entropy in the hyperposterior $P_{t}$ is shown over time. Across the three experiments, the hyperposterior of HP-GP-TS concentrates more than that of MAP GP-TS - which could be due to the greedy prior selection. As a comparison, HP-GP-TS places more than 95% of the probability mass on one prior (based on the reference values) in the lengthscale experiment where as the PE-methods barely removed priors.
@@
Figure 7: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ across all time steps of the kernel experiment.
@@
In Fig. 7, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the synthetic experiments as confusion matrices. PE-GP-UCB almost always selects the prior using the Matérn-3/2 kernel. This prior induces a distribution over functions that are less smooth compared to the other kernels. Hence, the Matérn-3/2 kernel produces much higher confidence intervals outside the observed data and is thus more optimistic, leading to excessive exploration. PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed. The overall “accuracy” of the selected priors, i.e. ${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$ where $\mathds{1}$ is the indicator function, of the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively. For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels. However, the RBF, Matérn and RQ kernels are often confused with each other. These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors given enough data.
@@
Figure 8: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to ±$\pm 1$ standard error.
@@
We consider two setups with real-world data. Using the Intel Berkeley dataset Madden et al. (2004), the first setup considers 46 temperature sensors with measurements over 19 days. The goal is to select the sensor with the highest temperature. The data is split into 12 training days and 7 test days. For each day in the training set, the empirical mean and covariance is computed and set as a prior. For each run, the unknown function $f$ is selected as a set of measurements from the test data and the noise variance is set to $\sigma^{2}=0.7^{2}$.
The second setup uses sensor data from 211 sensors along the I-880 highway in March 2023 from the California Performance Measurement System (PEMS) Chen et al. (2001); California Department of Transportation (2024). The goal is to select the sensor with the lowest speed in order to identify congestions. The sensor data prior to March 21st is used as training data and the rest is used as test data. Only sensor readings between 6AM and 13AM are used. For each hour in the training set (across all training days), the empirical mean and covariance is again used to construct a prior - yielding 7 priors. Again, the test data is used to sample the unknown function $f$ and the noise variance is set to $\sigma^{2}=2.25^{2}$. PE-GP-UCB and MAP-GP-TS are used as baselines to compare against PE-GP-TS and HP-GP-TS. The elimination methods use $\delta=0.05$, the horizon is set to $T=500$ and all methods are evaluated on $500$ seeds [^1] One outlier was removed from the Intel experiment, see Appendix B for more details..
@@
The cumulative regret for the real-world data experiments is presented in Fig. 8. For the Intel temperature data, HP-GP-TS and MAP-GP-TS obtain the lowest and second lowest cumulative regret respectively. For the PeMS data, PE-GP-TS obtains the lowest cumulative regret of all methods. HP-GP-TS and MAP-GP-TS have lower regret than PE-GP-UCB but their slopes are almost linear for the final time steps. To understand this better, we visualize quantiles of the total regret in Fig. 9 and the median cumulative regret in Fig. 11 in Appendix B. The figures show that MAP- and HP-GP-TS have the lowest median regret for both experiments and hence perform best in a majority of instances. However, the 90th and 95th quantiles are considerably larger for the PeMS data which impacts the average regret significantly. Hence, for the PeMS data, the prior elimination methods seem to yield more stable results.
The two right-most columns of Figs. 5 and 6 shows the average amount of active priors and the average entropy over time in the Intel and PeMS experiment. For the PeMS data, barely any priors are eliminated whilst for the Intel data around 0.75 to 0.80 priors are eliminated on average for the PE-methods. For the Intel data, the HP- and MAP GP-TS place around than 80% to 90% of the posterior probability mass in one prior based on the reference values. On the PeMS data, the posteriors concentrate more heavily - which could be premature given the average regret we observed.
(a) Intel
(b) PeMS
Figure 9: Quantiles of the final cumulative regret on the real-world data experiments. The median is highlighted in green. The whiskers correspond to the 5th and 95th percentile and the lower and upper edge of the box show the first and third quantile.
6 Discussion
Limitations
The main limitation of all the methods we have considered is that they require the set of priors $P$ to be discrete. In addition, their computational cost scales linearly with the number of priors considered. In theory, the elimination-based methods could have lower computational cost compared to HP-GP-TS. However, in practice, priors are rarely eliminated as shown by Fig. 5. At best, one sixth of the priors is eliminated in the kernel experiment for PE-GP-TS. Since only priors that are selected can be eliminated and the confidence parameters $\beta_{t}$, $\xi_{t}$ increasing over time, including more priors could lead to less priors being eliminated overall.
Future work
To lower the computational complexity of the elimination methods, one would either need tighter confidence bounds for the elimination criteria or need to allow more than the selected priors to be eliminated. The bi-level sampling approach of HP-GP-TS provides efficient exploration but suffers from higher computational cost since it cannot discard priors. An interesting line of future work could study how the probabilities in the hyperposterior could be used to discard priors that are unlikely to be selected in the remaining rounds. It would also be interesting to study selection criteria for the prior $p_{t}$ that maximize the disagreement among the priors such that evidence for the true prior is generated quicker.
A key difference between PE-GP-TS and the meta bandits methods of Kveton et al. (2021); Basu et al. (2021); Hong et al. (2022) is that they consider multiple bandit instances sampled from the true prior $p^{*}$. Specifically, Hong et al. (2022) study a setting with arbitrary interleaving of the instances. Generalizing our results for HP-GP-TS to a similar setting would provide an interesting complement to these methods beyond the standard and linear setting.
@@ 7 Conclusion
In this paper, we have proposed two algorithms for joint prior selection and regret minimization in GP bandits based on GP-TS. We have analyzed the algorithms theoretically and provided regret bounds whose order closely match previous work. We have experimentally evaluated both algorithms on both synthetic and real-world data. We find that they both select the true prior more often and obtain lower regret than previous work due to lowering the amount of optimistic exploration.
Acknowledgements
@@ A.1 PE-GP-TS
Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$. Additionally, for Eq. 5 we must take an additional union bound over $p\in P$.
@@
Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$y taking the union bound over ${\mathcal{X}}$, $P$ and $t\geq 1$, Eq. 5 holds w.p. at least $21-\delta/$ 2. By the same reasoning and skipping the union bound over $P$, Eq. 4 holds w.p. at least $21-\delta/$ 2. Thus, both events hold w.p. at least $1-\delta$. ∎
Next, we state three lemmas from Ziomek et al. (2024b) that are used in the proof of our regret bound.
@@
(Lemma 5.1 of Ziomek et al. (2024b)) If $\xi_{t}=2\sigma^{2}\log\left(\frac{|P|\pi^{2}t^{2}}{6\delta}\right)$, then the following holds with probability at least $1-\delta$:
@@
(Lemma 5.2 of Ziomek et al. (2024b)) Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$, then if $\mu_{1,p^{*}}$ and $k_{1,p^{*}}$ satisfy $|\mu_{1,p^{*}}(\cdot)|<\infty$ and $k_{1,p^{*}}(\cdot,\cdot)\leq 1$ and Section 4 holds, then
@@
(Lemma 5.3 of Ziomek et al. (2024b)) For $C=2/\log(1+\sigma^{-2})$, $\sum_{t\notin{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\leq\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$ where $\beta_{T}=\max_{p\in P}\beta_{T}$ and $\hat{\gamma}_{T}=\max_{p\in P}{\gamma}_{T,p}$.
@@
Then, we lower bound $f(x_{t})$
[…]
Combining, Eqs. 18 and 20 we obtain
@@
First, we show that the true prior $p^{*}$ is never rejected if Sections 4 and A.1 hold.
@@
[…]
@@
Using Eqs. 14 and 21, we can bound the cumulative regret as follows:
@@
where $B_{p^{*}}:=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$. If $t\notin{\mathcal{C}}$, line 9 in Algorithm 1 evaluates to false and hence
@@
Additionally, using Section A.1, we can bound the Gaussian noise:
@@
Finally, applying Eqs. 15 and 4.1, we obtain the result
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\gamma_{T,p^{*}}}+2\sqrt{CT\beta_{T}\gamma_{T}|P|}.$ (33)
A.2 HP-GP-TS
First, we state and prove Theorem 4.1. See 4.1
Proof.
First, note that $I_{p}({\mathbf{y}}_{T};f)=H({\mathbf{y}}_{T}|p)-H({\mathbf{y}}_{T}|f,p)$. As in Srinivas et al. (2012), $H({\mathbf{y}}_{T}|f,p)$ is simply the entropy of $T$ zero-mean Gaussians with variance $\sigma^{2}$: $\frac{T}{2}\log(2\pi e\sigma^{2})$. Using the chain rule of entropy, $H({\mathbf{y}}_{T}|p)=H({\mathbf{y}}_{T-1}|p)+H(y_{T}|{\mathbf{y}}_{T-1},p)=H({\mathbf{y}}_{T-1}|p)+\frac{1}{2}\log(2\pi e(\sigma^{2}+\sigma^{2}_{t}(x_{T}))$ where the second equality is due to $y_{T}|{\mathbf{y}}_{T-1},p\sim{\mathcal{N}}(\mu_{T,p}(x_{T}),\sigma^{2}_{T,p}(x_{T})+\sigma^{2})$. The rest follows by induction.
The second part follows from the proof of Lemma 5.4 in Srinivas et al. (2012):
$\sigma_{t,p}^{2}(x_{t})$ $=\sigma^{2}(\sigma^{-2}\sigma^{2}_{t,p}(x_{t}))$ (34)
$\leq\sigma^{2}\frac{\sigma^{-2}}{\log(1+\sigma^{-2})}\log(1+\sigma^{-2}\sigma^{2}_{t,p}(x_{t}))$ (35)
$\leq\frac{1}{\log(1+\sigma^{-2})}\log(1+\sigma^{-2}\sigma^{2}_{t,p}(x_{t})).$ (36)
Using Eq. 8, we get the desired result. ∎
Based on Theorem 4.1, Theorem 4.1 follows almost immediately.
See 4.1
Proof.
From Theorem 4.1, we have for any fixed sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and prior $p^{*}\in P$, that ${\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}\leq CI_{p^{*}}({\mathbf{y}}_{T};f)$. The mutual information $I_{p^{*}}({\mathbf{y}}_{T};f)$ given the prior $p^{*}$ is in turn bounded by the maximal information gain $\gamma_{T,p^{*}}$ for the same prior. Hence,
${\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\right]\leq{\mathbb{E}}\left[C\gamma_{T,p^{*}}\right]={\mathbb{E}}_{p^{*}\sim P_{0}}\left[C\gamma_{T,p^{*}}\right].$ (37)
Then, we are ready to state and prove our regret bound for HP-GP-TS.

### v2

@@ 5 Experiments
We consider three synthetic setups with different choices of priors in $P$. For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$. All kernels use a lengthscale of $1.0$ and are scaled s.t. $k(x,\tilde{x})\leq 1$. In addition, the mean function for all priors is zero everywhere. For the second setup, the priors use the RBF kernel with lengthscales $4$, 22, 11 or $1/2$. For the third setup, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$. Dimensions larger than $5$ are wrapped around 1, i.e. $((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize. All priors use the RBF kernel with lengthscale $\ell=8$. For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$. For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$. The prior elimination methods use $\delta=0.05$. All models are evaluated on 500 seeds on each setup. As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation.. Regardless of the selected prior, the weighting of the hyperposterior $P_{t}$ is updated for all priors. Hence, greedily selecting the prior could reduce unnecessary exploration. In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{1}=\{p^{*}\}$.
@@
The cumulative regret for the three synthetic experiments is shown in Fig.˜3. Across all three experiments, we observe that HP-GP-TS has lower regret than the other methods and performs close to the oracle GP-TS. For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB. Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods. MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment. The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances.
The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Fig.˜4. The PE-methods eliminate at most one prior on average. In contrast, the hyperposterior entropy of HP-GP-TS is equivalent to 80-90% of the probability mass being assigned to one prior. HP- and MAP-GP-TS thus effectively discards priors at a much faster rate. The same pattern holds for the lengthscale and subspace experiments, see Figs.˜9 and 10 in Appendix˜D.
@@
Figure 5: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ for the kernel experiment. Row-wise normalized to 100%.
@@
In Fig.˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices. PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds. The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration. PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed. The overall “accuracy" of the selected priors, i.e. ${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively. For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels. However, the RBF, Matérn and RQ kernels are often confused with each other. These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data. See Fig.˜11 in Appendix˜D for confusion matrices in the lengthscale and subspace experiments.
Figure 6: Total regret for the lengthscale and subspace experiments as the number of priors $|P|$ increases.
Scaling $|P|$
We perform two experiments to understand how the regret of our algorithms scale with the number of priors. In both experiments, the average difficulty of the problem is kept constant such that the regret of the oracle models is constant. In the first experiment, we increase the discretization of the lengthscale values. The lengthscales are equidistantly spaced in $[0.5,4]$ with $|P|\in\{8,16,32,64,128\}$. As $|P|$ increases, the difference between similar priors is reduced. In the second experiment, we increase the number of priors in the subspace experiment from 5 up to 16. Each prior can share at most 3 out of 4 dimensions with other priors which ensures the priors remain meaningfully different. The total regret as the number of priors increases is shown in Fig.˜6. For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm. This is likely due to the redundancy in the priors. However in the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle.
@@
Figure 7: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to $\pm 1$ standard error. The average final regret for PE-GP-UCB is 511.9 for PNW.
@@
We perform three experiments with real-world data from the Intel Berkeley dataset (Madden et al., 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann & Bretherton, 1999, 2000). Each dataset contains measurements from a set of sensors over time. We split each dataset into a training and test set where the test set contains the last third of the data. Hence, the distribution of the test data may have shifted slightly from the training data. The training sets are split further into separate buckets to define our priors. For each bucket $p$, we compute the empirical mean $\hat{\mu}_{p}$ and covariance $\hat{\Sigma}_{p}$ which defines the prior ${\mathcal{GP}}(\hat{\mu}_{p},\hat{\Sigma}_{p})$. The buckets in the Intel data corresponds to the 12 days in the training dataset. For the PeMS data, each hour between 06:00 and 13:00 defines one prior, giving 7 priors. For the daily precipication data, each month in the year constitutes a prior, yielding 12 priors. When running the experiments, we select a measurement of all sensors from the test data uniformly at random. The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipication respectively for the three datasets. When the algorithms select an arm to evaluate, we add Gaussian noise to $y_{t}$ with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al. (2012); Bogunovic et al. (2016). See Appendix˜C for more details about the experimental setup.
@@
The cumulative regret for the experiments with real-world data is presented in Fig.˜7. For the Intel and PNW data, HP- and MAP GP-TS obtain the lowest and second lowest cumulative regret respectively. HP- and MAP GP-TS have lower regret than PE-GP-TS initially but PE-GP-TS catches up and has the lowest total regret. To understand this better, we visualize quantiles of the total regret in Fig.˜13. MAP- and HP-GP-TS have the lowest median regret for all three experiments and hence perform best in a majority of instances. However, the 90th and 95th quantiles are considerably larger for the PeMS data which impacts the average regret significantly. Hence, for the PeMS data, the prior elimination methods seem to yield more stable results.
In Fig.˜4, the number of priors remaining in $|P_{t}|$ and the hyperposterior entropy is shown for the Intel experiment. Similar to the synthetic experiment, on average, the prior elimination methods eliminate less than 1 prior whereas the hyperposteriors of HP- and MAP GP-TS concentrate the equivalent of 80-90% of the probability mass to one prior. The results for PeMS and PNW experiment are shown in Figs.˜9 and 10 in Appendix˜D. Here, effectively no priors are eliminated. For the PNW experiment, the hyperposteriors do not concentrate as much compared to the other experiments. This could indicate that knowing the exact prior is not as important for the PNW data.
@@ 6 Conclusion
In this paper, we have proposed two algorithms for joint prior selection and regret minimization in GP bandits based on GP-TS. We have analyzed the algorithms theoretically and have experimentally evaluated both algorithms on synthetic and real-world data. We find that they both select the true prior more often and obtain lower regret than previous work due to lowering the amount of optimistic exploration.
Acknowledgments
@@ A.1 PE-GP-TS
Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$. Additionally, for Eq.˜5 we must take an additional union bound over $p\in P$.
@@
Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Eq.˜5 holds w.p. at least $1-\delta/2$. By the same reasoning and skipping the union bound over $P$, Eq.˜4 holds w.p. at least $1-\delta/2$. Thus, both events hold w.p. at least $1-\delta$. ∎
Next, we state three lemmas from Ziomek et al. (2025) that are used in the proof of our regret bound.
@@
(Lemma 5.1 of Ziomek et al. (2025)) If $\xi_{t}=2\sigma^{2}\log\left(\frac{|P|\pi^{2}t^{2}}{6\delta}\right)$, then the following holds with probability at least $1-\delta$:
@@
(Lemma 5.2 of Ziomek et al. (2025)) Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$, then if $\mu_{1,p^{*}}$ and $k_{1,p^{*}}$ satisfy $|\mu_{1,p^{*}}(\cdot)|<\infty$ and $k_{1,p^{*}}(\cdot,\cdot)\leq 1$ and Section˜4.1 holds, then
@@
(Lemma 5.3 of Ziomek et al. (2025)) For $C=2/\log(1+\sigma^{-2})$, $\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\leq\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$ where $\beta_{T}=\max_{p\in P}\beta_{T}$ and $\hat{\gamma}_{T}=\max_{p\in P}{\gamma}_{T,p}$.
@@
$\leq\tilde{f}_{t,p_{t}}(x_{t})+2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*}).$ $(\text{TS selection rule})$ (18)
Then, we lower bound $f(x_{t})$
$f(x_{t})$ $=\mu_{t,p_{t}}(x_{t})+\eta_{t}-\epsilon_{t}$ $(\text{Def. of }\eta_{t})$ (19)
@@
Combining, Eqs.˜18 and 20 we obtain
@@
First, we show that the true prior $p^{*}$ is never rejected if Sections˜4.1 and A.1 hold.
@@
$\leq\left|\sum_{i\in S_{t,p^{*}}}\epsilon_{i}\right|+\sum_{i\in S_{t,p^{*}}}\left|f(x_{i})-\mu_{i,p^{*}}(x_{i})\right|$ $(\text{Triangle ineq.)}$ (23)
@@
Using Eqs.˜14 and 21, we can bound the cumulative regret as follows:
@@
where $B_{p^{*}}:=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$. If $t\notin\mathcal{C}$, line˜9 in Algorithm˜1 evaluates to false and hence
@@
Additionally, using Section˜A.1, we can bound the Gaussian noise:
@@
Finally, applying Eq.˜15, we obtain the result
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (33)
A.2 UCB-analysis of HP-GP-TS
Next, we state and prove our UCB-based regret bound for HP-GP-TS.